Find the number that logically continues each of these series:
a) 2,3,5,9,17,
b) 14,19,29,40,44,52,

Answer by checkley71(8403)

2,3,5,9,17,33 [FACTORS ARE ADD 1,2,4,8,16]
14,19,29,40,44,52,59 [ADD 1+4+14=19, ADD 1+9+19=29, ADD 2+9+29=40, ETC.]